A sample of n=18 observations is drawn from a normal population with mean=960 and standard deviation= 220. Find each of the following:

A. P(X>1037)

B. P(C<877)

C. P(X>934)


1
Expert's answer
2021-10-17T16:37:02-0400

n = 18

µ = 960

σ = 220

z = (x – µ)/ (σ/"\sqrt{n}" )

(a)   z = (1037 – 960)/(220/"\sqrt{18}" ) = 1.48

Using technology, P(z > 1.48) = 0.0688

(b)  z = (877 – 960)/(220/"\sqrt{18}" ) = –1.6

Using technology, P(z < –1.6) = 0.0547

(c)   z = (934 – 960)/(220/"\sqrt{18}" ) = –0.5

Using technology, P(z > –0.5) = 0.692